A trader buys some eggs at 6naira each. She finds out that six of them are broken, she sells the rest at 10naira each and makes a profit of 160. How many eggs did she buy?
Answer by ikleyn(52797)

.
A trader buys some eggs at 6naira each. She finds out that six of them are broken,
she sells the rest at 10naira each and makes a profit of 160. How many eggs did she buy?
~~~~~~~~~~~~~~


Let n be the number of eggs she bought.


Then we have this equation


    Profit = Revenue - Cost;

    Profit = 160;  Cost = 6n;  Revenue = 10*(n-6);


       +---------------------------+
       |    160 = 10*(n-6) - 6n.   |
       +---------------------------+


Simplify it


    160 = 10n - 60 - 6n

    160 + 60 = 4n

      220    = 4n

       n     = 220/4 = 55.


ANSWER.  She bought 55 eggs.


CHECK,   10*(55-6) = 10*49 = 490;   6n = 6*55 = 330;   490 - 330 = 160  (profit).   ! Correct !

Solved.